OpenStudy (anonymous):

PLEASE HELP!!! Solve 4^x + 7 = 6^x - 1 a.) x ≈ 0.04 b.) x ≈ 0.05 c.) x ≈ 19.51 d.) x ≈ 28.35

OpenStudy (anonymous):

i am going to make a guess my guess is that this is really \[4^{x+7}=6^{x-1}\] am i right?

OpenStudy (anonymous):

yes your right!

OpenStudy (anonymous):

start with \((x+7)\ln(4)=(x-1)\ln(6)\) and solve using a raft of algebra

OpenStudy (anonymous):

i already did this problem i got 0.04 as my answer was i correct?

OpenStudy (anonymous):

what did you get before you used a calculator?

OpenStudy (anonymous):

and no, it is not correct